OpenStudy (anonymous):

Solve the equation cos x° = sin 40°, where 0 < x < 90.

OpenStudy (callisto):

For 0 < x < 90, cos (90-x) = sin (x), and sin (90-x) = cos(x)

OpenStudy (anonymous):

I still dont know how to get the answer lol

OpenStudy (callisto):

Ehm.... Pick one identity from above first.

OpenStudy (anonymous):

cos (90-x) = sin (x), is that what you mean?

OpenStudy (callisto):

You can pick either one lol So, your choice is cos(90-x) = sin(x). Let's just rename x as "a". So, we have cos(90-a) = sin(a) Compare this with what you have now: sin 40 What is a in this case?

OpenStudy (anonymous):

40? I dont know I'm terrible at geometry

OpenStudy (callisto):

40 is right, and it's not geometry, haha! Now, substitute a by 40, you have cos(90-50) = sin(40). Can you simplify the left side?

OpenStudy (anonymous):

Oh well this a question on my geometry test lol. 90-50 is 40

OpenStudy (callisto):

Aiya! My bad! It should be cos(90-40) = sin(40). Can you simplify the left side again?

OpenStudy (anonymous):

90-40=50

OpenStudy (callisto):

Yup! So, you have cos(50) = sin(40). I guess that's all?!

OpenStudy (anonymous):

Yes, thanks a lot!

OpenStudy (callisto):

You're welcome~